### v2

# Non-perturbative renormalization group for Higgs-like models in 4D

Abstract
We recently defined a model of 2 coupled SU(2) Higgs-like doublets which revealed an interesting structure of renormalization group flows to 1-loop when the SU(2) is broken to U(1). In this article we compute the beta functions to 3 loops and show that the 1-loop structure of flows persists to higher orders. For SU(2) broken to U(1), we conjecture a beta function to all orders. The flows can be extended to large coupling using a strong-weak coupling duality $g\to 1/g$. One finds a line of fixed points which are new conformal field theories in 4 spacetime dimensions which are non-unitary due to negative norm states but still have real eigenvalues. We also find massless flows between 2 non-trivial fixed points, and a regime with a cyclic RG flow. For the flows between points on the critical line, we compute the anomalous dimensions of the perturbations in the UV and IR, and identify some special points where anomalous dimensions are rational numbers. The model is non-unitary since the hamiltonian is pseudo-hermitian, $H^{\dagger}={\cal K}H{\cal K}^{\dagger}$. The unitary operator ${\cal K}$ satisfies ${\cal K}^{2}=1$ and this allows a projection onto positive norm states with a unitary time evolution with positive probabilities.
I Introduction
There is a conspicuous shortage of well-understood conformal field theories (CFT’s) in 4 spacetime dimensions that aren’t simply free fields, as for asymptotically free QCD Wilczek ; Gross . A well known exception is ${\cal N}=4$ SUSY Yang-Mills theory. For marginal perturbations, such as standard generalizations of $\phi^{4}$ for the Higgs potential, the renormalization group (RG) flows are rather limited in scope in exactly 4 spacetime dimensions. For scalar fields, with $\phi^{4}$ interactions, using the epsilon expansion in $D=4-\epsilon$ dimensions, one can find CFT’s in 3 spacetime dimensions such as the Wilson-Fisher fixed points, which have irrational anomalous dimensions WilsonFisher . In fact, this is at the heart of the so-called hierarchy problem, namely that straightforward RG arguments would indicate a much higher mass for the Higgs boson since there is no known ultra-violet (UV) fixed point. [^2] Without a known UV completion of the Higgs sector, there may even be higher dimension operators that make the theory non-unitary. This article will eventually scrutinize the idea that all quantum field theories (QFT’s) begin and end at RG fixed point CFT’s, at least for non-unitary theories.
By contrast, in 2D one has a large toolkit of algebraic structures to study CFT’s, in particular the Virasoro algebra and affine Lie algebra symmetry of WZW models BPZ ; WittenWZW ; KZ . For the purposes of this article, we should point out that in the classification of minimal models of CFT based on the Virasoro algebra, one finds many more non-unitary theories than unitary, all with real eigenvalues of the hamiltonian, and many of them have important physical applications, such as the Lee-Yang edge singularity. Here the non-unitarity is manifested in negative norm states, rather than non-hermiticity of the hamiltonian. We would also like to stress that constructions of these CFT’s often requires projecting onto unitary sub-Hilbert spaces of non-unitary theories FQS . For instance, in the Coulomb gas method one must project the Hilbert space of a free boson to the proper space of the minimal model CFT DotsenkoFateev , and this can be described as a BRST procedure BernardFelder , or a reduction based on the SU(2) quantum group. In this way one can describe both the unitary and non-unitary minimal models of CFT in 2D. This procedure is similar to the BRST treatment of gauge theories which consistently projects out ghosts. We should also mention that the unitary CFT’s can be understood as coset reductions of current-algebras GKO , so that current algebras play a fundamental role in the classification of CFT’s.
It would certainly be nice to have such a 2D playground in 4D, and we hope this article opens such a door, however narrow it may turn out to be, based on the model we recently defined ALDolls . There we constructed some new kinds of marginal perturbations of scalar fields in 4D that have a rich structure of RG flows, at least to 1-loop. The original goal of that work was to define Higgs-like models with more interesting RG flows than the usual ones. A line of ultraviolet (UV) or infrared (IR) fixed points, and perhaps more interestingly, some cyclic RG flows were found, but based only on the 1-loop approximation. [^3] The possibility of cyclic RG flows was considered long ago in Wilson’s pioneering work on the RG KWilson . The nature of these marginal operators are such that they satisfy an interesting operator product expansion with an SU(2) Lie-algebraic structure, similar to current algebras in 2D KZ . These models can be defined for an arbitrary Lie algebra, however with applications to Higgs physics in mind, we specialized to ${\rm SU(2)}$. This algebraic structure leads to simple formulas for the 1-loop beta functions. Since the cyclic RG flows require flowing to infinite values of the couplings, it is necessary to check if the 1-loop behavior persists to higher orders. This is main subject of this paper, which can be viewed as Part II of the RG aspects of our previous work ALDolls . We will show that the 1-loop features do indeed persist non-perturbatively in a very interesting, non-trivial manner.
The regime of couplings where SU(2) is broken to U(1) is our main interest, which has 2 independent couplings. It turns out that one direction in the 2 coupling constants of our model corresponds to a line of fixed points, which are new CFT’s in 4D which we develop further below. In ALDolls we studied spontaneous symmetry breaking of SU(2) to U(1) and speculated that the cyclic regime could potentially resolve the hierarchy problem for Higgs physics. Furthermore, we found that the model has an infinite number of vacuum expectation values with Russian Doll scaling. We speculated that this could potentially explain the origin of families in the Standard Model, wherein each family is one of an infinite number of Russian Dolls which appear as resonances in scattering. The interplay between the Higgs mechanism and cyclic RG flows is clearly something worth developing further than in ALDolls , however in this article we leave this more speculative issue aside and focus solely on providing firm foundations for the quantum field theories involved based on their algebraic structures, and study their fixed points and RG flows, which are interesting in and of themselves.
As defined in ALDolls and reviewed below, the hamiltonian $H$ for our models is not hermitian, but pseudo-hermitian, which makes the model non-unitary. This is not an issue for applications to statistical mechanics in 4 spatial dimensions. For quantum mechanics in $3+1$ spacetime dimensions one needs additional positivity constraints related to unitarity. This doesn’t necessarily entail an unruly hornets’ nest if the hamiltonian has some additional algebraic structure. On the other hand, if one has merely the statement $H^{\dagger}\neq H$, then there is not even a nest to speak of. Let us mention that there has recently been renewed and growing interest in non-unitary minimal models in 2D and RG flows between them MussardoLG ; Tanaka ; KlebanovMM ; Negro . In BL38 hamiltonians satisfying $H^{\dagger}={\cal K}H{\cal K}^{\dagger}$, with ${\cal K}$ unitary, were classified according to ${\cal C},{\cal P},{\cal T}$ discrete symmetries, and 38 universality classes of random hamiltonians were obtained, extending the 3-fold classifications of Wigner and Dyson based on time-reversal symmetry, and the 10-fold classifications of Altland-Zirnbauer which included particle-hole symmetry Dyson ; AltlandZirnbauer . These classes have seen many applications to open quantum systems. The models considered in this paper are a special case where
$H^{\dagger}={\cal K}\,H\,{\cal K}^{\dagger},~{}~{}~{}{\cal K}^{\dagger}{\cal K}=1,~{}~{}~{}~{}~{}{\cal K}^{\dagger}={\cal K},~{}~{}~{}~{}\Longrightarrow~{}~{}~{}{\cal K}^{2}=1.$ (1)
We will argue below that this structure leads to a consistent unitary quantum mechanics if one uses the operator ${\cal K}$ to project onto positive norm states. In fact pseudo-hermitian hamiltonians were proposed as consistent extensions of quantum mechanics long ago by Pauli Pauli . More recently pseudo-hermitian quantum mechanics has been developed in detail by Mostafazadeh and others Mosta1 ; Mosta2 ; Mosta3 , and is by now a well-established and highly cited framework, although one still expects some resistance. Pseudo-hermiticity as defined by (1) is not the same as ${\cal P}{\cal T}$ symmetric quantum mechanics, which also has real eigenvalues if the ${\cal P}{\cal T}$ symmetry is unbroken BenderReview .
Let us summarize the main results and organization of this article. In the next section we review the definition of the models in ALDolls . In Section III we address the non-unitarity issue and propose how to project onto positive norm states which leads to a consistent quantum mechanics, with a unitary time evolution of positive probabilities. In Section IV we compute the RG beta functions to 3-loops using just the OPE structure for fully anisotropic SU(2) perturbations, which has 3 independent couplings. In Section V we specialize to SU(2) broken to U(1), which involves only 2 couplings. Based on the structure of the 3-loop results we propose beta functions to all orders. In Section VI we study the RG flows in the resulting 2-coupling parameter space to all orders in the couplings. This is made possible by a certain RG invariant $Q$ where RG flow trajectories are constant $Q$ contours, and also a strong-weak coupling duality $g\to 1/g$ of the beta functions. The resulting picture of RG flows presents itself as a kind of master flowchart for 2 couplings, since it includes new fixed point CFT’s, marginally relevant or irrelevant, asymptotic freedom, massless flows between non-trivial fixed points, and alas cyclic RG flows. [^4] This master flowchart is natural if one starts with a single coupling, which has a marginally relevant and irrelevant direction. Opening up this flow with an additional coupling that preserves this flow along the diagonals leads to this kind of flowchart. We refer to the figures below. For the cyclic flows, the RG period $\lambda$ is a simple function of the RG invariant $Q$, namely $\lambda=2\pi/\sqrt{Q}$. The existence of a regime of couplings where the RG is cyclic calls into question the commonly accepted paradigm that every QFT begins or ends at a fixed point, and we comment on this in that section. By taking one of the couplings to be imaginary, we also find massless flows between different fixed points along the critical line. The latter are 4D versions of flows between non-unitary minimal models in 2D which have received much attention recently MussardoLG ; Tanaka ; KlebanovMM ; Negro .
II Definition of the models
Our models are defined as certain marginal operator perturbations of a conformal field theory consisting of two free Higgs doublets. Introduce two independent ${\rm SU(2)}$ doublets of complex bosonic spin-0 fields, $\Phi_{i}$ and $\widetilde{\Phi}_{i}$, ${i=1,2}$ and their usual hermitian conjugates $\Phi^{\dagger}$, $\widetilde{\Phi}^{\dagger}$. Under ${\rm SU(2)}$ transformations, $\Phi(x)\to U\Phi(x)$, where $U$ is a $2\times 2$ unitary ${\rm SU(2)}$ group matrix acting on the “i” indices in the above equation and the same for $\widetilde{\Phi}$. The symmetry is actually $U(2)$ which allows an additional U(1) “hypercharge” however this will not be essential to this article. The free action is the standard one for free massless bosonic fields.
${\cal S}_{0}=\int d^{4}x\left(\partial_{\mu}\Phi^{\dagger}\partial^{\mu}\Phi+\partial_{\mu}\widetilde{\Phi}^{\dagger}\partial^{\mu}\widetilde{\Phi}\right),$ (2)
where $\Phi^{\dagger}\Phi=\sum_{i=1,2}\Phi^{\dagger}_{i}\Phi_{i}$ etc. The free theory has global ${\rm SU(2)}\otimes{\rm SU(2)}$ symmetry since the fields $\Phi,\widetilde{\Phi}$ are independent, however the interactions introduced below will break this down to the diagonal ${\rm SU(2)}$. The quantization of this theory is standard. Consider the quantization of one component $\Phi=\Phi_{i}$ for $i=1~{}{\rm or}~{}2$, since the same applies to $\widetilde{\Phi}$. Expand the field in terms of operators $a_{\pm},a^{\dagger}_{\pm}$:
$\Phi(x)$ $=$ $\int\dfrac{d^{3}{\bf{k}}}{(2\pi)^{3/2}}\dfrac{1}{\sqrt{2\omega_{\bf{k}}}}\left(a^{\dagger}_{-}({\bf{k}})e^{-ik\cdot x}+a_{+}({\bf{k}})e^{ik\cdot x}\right)$
$\Phi^{\dagger}(x)$ $=$ $\int\dfrac{d^{3}{\bf{k}}}{(2\pi)^{3/2}}\dfrac{1}{\sqrt{2\omega_{\bf{k}}}}\left(a_{-}({\bf{k}})e^{ik\cdot x}+a^{\dagger}_{+}({\bf{k}})e^{-ik\cdot x}\right),$ (3)
where $k\cdot x=\omega_{\bf{k}}t-{\bf{k}}\cdot{\bf{x}}$ with $\omega_{\bf{k}}=|{\bf{k}}|$, and similarly for $\widetilde{\Phi}$. Canonical quantization of the boson field yields
$[a_{\pm}({\bf{k}}),a^{\dagger}_{\pm}({\bf{k}}^{\prime})]=\delta^{(3)}({\bf{k}}-{\bf{k}}^{\prime}),~{}~{}~{}~{}[a_{\pm}({\bf{k}}),a^{\dagger}_{\mp}({\bf{k}}^{\prime})]=0,$ (4)
and the hamiltonian is
$H_{0}=\int d^{3}{\bf{k}}\,\omega_{{\bf{k}}}\left(a^{\dagger}_{+}({\bf{k}})\,a_{+}({\bf{k}})+a^{\dagger}_{-}({\bf{k}})\,a_{-}({\bf{k}})\right).$ (5)
In defining the marginal perturbations of interest, it’s necessary to introduce a discrete symmetry operator ${\cal K}$, satisfying
${\cal K}a_{\pm}({\bf{k}}){\cal K}=\pm a_{\pm}({\bf{k}}),~{}~{}~{}{\cal K}a^{\dagger}_{\pm}({\bf{k}}){\cal K}=\pm a^{\dagger}_{\pm}({\bf{k}}).$ (6)
The operator ${\cal K}$ is unitary:
${\cal K}^{\dagger}{\cal K}=1,~{}~{}~{}~{}{\cal K}={\cal K}^{\dagger},~{}~{}~{}~{}\Longrightarrow~{}~{}~{}{\cal K}^{2}=1.$ (7)
Then the free hamiltonian is ${\cal K}$ invariant:
${\cal K}H_{0}{\cal K}=H_{0}=H^{\dagger}_{0}.$ (8)
Consider standard parity ${\cal P}$ and time-reversal symmetry ${\cal T}$. Since parity flips the sign of momentum:
${\cal P}a_{\pm}({\bf{k}}){\cal P}=a_{\pm}(-{\bf{k}}),~{}~{}~{}{\cal P}a^{\dagger}_{\pm}({\bf{k}}){\cal P}=a^{\dagger}_{\pm}(-{\bf{k}}),~{}~{}~{}~{}~{}~{}\Longrightarrow~{}~{}~{}~{}{\cal P}\Phi(t,{\bf{x}}){\cal P}=\Phi(t,-{\bf{x}}),~{}~{}~{}{\cal P}H_{0}{\cal P}=H_{0}.$ (9)
Time-reversal also flips the sign of ${\bf{k}}$ and only differs from ${\cal P}$ in that it is anti-unitary, namely ${\cal T}z{\cal T}=z^{*}$ for $z$ a complex number. This leads to
${\cal T}a_{\pm}({\bf{k}}){\cal T}=a_{\pm}(-{\bf{k}}),~{}~{}~{}{\cal T}a^{\dagger}_{\pm}({\bf{k}}){\cal T}=a^{\dagger}_{\pm}(-{\bf{k}})~{}~{}~{}~{}\Longrightarrow~{}~{}~{}{\cal T}\Phi(t,{\bf{x}}){\cal T}=\Phi(-t,{\bf{x}}),~{}~{}~{}~{}~{}{\cal T}H_{0}{\cal T}=H_{0}.$ (10)
This implies
${\cal P}{\cal T}a_{\pm}({\bf{k}}){\cal P}{\cal T}=a_{\pm}({\bf{k}}).$ (11)
Comparing with (11), we can interpret the operators $a_{\pm}^{\dagger}$ as creating states that are even or odd under ${\cal P}{\cal T}$. It’s important to point out that unlike ${\cal P},{\cal T}$, the operator ${\cal K}$ does not act locally on fields.
Our models will involve $\Phi^{2}\widetilde{\Phi}^{2}$ marginal interactions involving the ${\cal K}$-conjugate fields $\Phi^{\dagger_{\scriptscriptstyle\kappa}}$ and $\widetilde{\Phi}^{\dagger_{\scriptscriptstyle\kappa}}$:
$\Phi^{\dagger_{\scriptscriptstyle\kappa}}(x)\equiv{\cal K}\,\Phi^{\dagger}{\cal K}=\int\dfrac{d^{3}{\bf{k}}}{(2\pi)^{3/2}}\dfrac{1}{\sqrt{2\omega_{\bf{k}}}}\left(-a_{-}({\bf{k}})e^{ik\cdot x}+a^{\dagger}_{+}({\bf{k}})e^{-ik\cdot x}\right).$ (12)
The above action of ${\cal K}$ on momentum space operators implies a non-local action on fields. Using $a_{\pm}|0\rangle=0$, due to the extra minus sign in the above equation one finds the free field 2-point correlation functions:
$\langle\Phi^{\dagger_{\scriptscriptstyle\kappa}}(x)\Phi(y)\rangle=-\langle\Phi(x)\Phi^{\dagger_{\scriptscriptstyle\kappa}}(y)\rangle=\dfrac{1}{4\pi^{2}|x-y|^{2}}.$ (13)
For any operator $A$ define it’s ${\cal K}$-conjugate as
$A^{\dagger_{\scriptscriptstyle\kappa}}\equiv{\cal K}\,A^{\dagger}{\cal K}.$ (14)
The operator $\Phi^{\dagger_{\scriptscriptstyle\kappa}}\Phi$ is not its own hermitian conjugate, but is invariant under ${\dagger_{\scriptscriptstyle\kappa}}$ conjugation:
$(\Phi^{\dagger_{\scriptscriptstyle\kappa}}\Phi)^{\dagger_{\scriptscriptstyle\kappa}}=\Phi^{\dagger_{\scriptscriptstyle\kappa}}\Phi.$ (15)
Motivated by this, define the operators
$J^{a}\equiv\Phi^{\dagger_{\scriptscriptstyle\kappa}}\sigma^{a}\Phi=\sum_{i,j}\Phi^{\dagger_{\scriptscriptstyle\kappa}}_{i}\sigma^{a}_{ij}\Phi_{j}~{}~{}~{}~{}~{}~{}(\sigma^{a})^{\dagger}=\sigma^{a},$ (16)
where $\sigma^{a}$ are the usual hermitian Pauli matrices:
$\sigma^{1}=\begin{pmatrix}0&1\\ 1&0\end{pmatrix},~{}~{}~{}\sigma^{2}=\begin{pmatrix}0&-i\\ i&0\end{pmatrix},~{}~{}~{}\sigma^{3}=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix},~{}~{}~{}{\rm Tr}(\sigma^{a}\sigma^{b})=2\delta^{ab}.$ (17)
and similarly for $\widetilde{J}^{a}$. The operators $J^{a}$, as defined this way, have an interesting operator product expansion (OPE):
$J^{a}(x)J^{b}(y)=-\dfrac{2\kappa\delta^{ab}}{16\pi^{4}|x-y|^{4}}-\dfrac{if^{abc}}{4\pi^{2}|x-y|^{2}}\,J^{c}(y)+\ldots~{}~{}~{}(\kappa=1),$ (18)
where the structure constants $f^{abc}$ are
$[\sigma^{a},\sigma^{b}]=if^{abc}\sigma^{c},~{}~{}~{}~{}~{}f^{abc}=2\epsilon^{abc},$ (19)
with $\epsilon^{abc}$ the completely anti-symmetric tensor with $\epsilon^{123}=1$. In the above equation $\kappa=1$, however we introduced $\kappa\neq 1$ for reasons that will become clear below. For ${\rm SU(2)}$ there is no sum over “$c$” in the equation (18) and this will simplify some calculations below. The extra minus sign for $\Phi^{\dagger_{\scriptscriptstyle\kappa}}$ in (12) is what leads to the commutator in (19) and the resulting structure constant $f^{abc}$ in the OPE (18). As we will see, the beta functions can be determined from the above OPE.
These are all the ingredients necessary to define our models. Define the marginal operators
${\cal O}^{a}(x)=J^{a}(x)\widetilde{J}^{a}(x),$ (20)
(no sum on $a$). The models are then defined by the action
${\cal S}={\cal S}_{0}+2\pi^{2}\int d^{4}x\,\sum_{a=1,2,3}g_{a}{\cal O}^{a}(x).$ (21)
The factor of $2\pi^{2}$ is introduced to simplify the beta-functions of the next section. When $g_{1}=g_{2}=g_{3}$ the model has SU(2) symmetry since it is built on the quadratic Casimir. The operators $J^{a}$ are not hermitian but pseudo-hermitian:
$J^{\dagger_{\scriptscriptstyle\kappa}}=J,~{}~{}~{}~{}~{}\widetilde{J}^{\dagger_{\scriptscriptstyle\kappa}}=\widetilde{J}.$ (22)
Thus the interacting hamiltonian is pseudo-hermitian:
$H^{\dagger}={\cal K}H{\cal K}~{}~{}~{}\Longrightarrow~{}~{}H^{\dagger_{\scriptscriptstyle\kappa}}=H.$ (23)
III Projecting onto positive norm states and unitarity
Our models could have applications to statistical mechanics in 4 spatial dimensions where the positivity constraints of unitarity are not so stringent. However, for potential applications to real Higgs physics in $3+1$ spacetime dimensions, a consistent quantum mechanics would be desirable, and this is the subject of this section. The subject of Quantum Mechanics for pseudo-hermitian hamiltonians is by now well developed, with extensive reviews Mosta1 ; Mosta2 ; Mosta3 . In this section we summarize the main properties as they specialize to our model.
(i)   Hamiltonian and Hilbert space
Assume we are given a Hilbert space ${\cal H}$ with a conventional positive definite inner product. Namely for $|\psi\rangle\in{\cal H}$, there exists an inner product such that $\langle\psi|\psi\rangle$ is positive definite. At time $t=0$ the states need to be properly normalized such that $\langle\psi|\psi\rangle=1$. There exists a basis in ${\cal H}$, $|\psi_{n}\rangle\in{\cal H}$ where for simplicity of notation we assume $n$ is discrete, $n=1,2,\ldots\infty$. Then
$|\psi\rangle=\sum_{n}c_{n}\,|\psi_{n}\rangle,~{}~{}~{}~{}\langle\psi_{m}|\psi_{n}\rangle=\delta_{nm},$ (24)
for some complex numbers $c_{n}$. Then
$\langle\psi|\psi\rangle=1,~{}~{}~{}~{}~{}~{}\Longrightarrow~{}~{}~{}\sum_{n}c_{n}^{*}c_{n}=1,$ (25)
where $c_{n}^{*}$ denotes complex conjugation. $|c_{n}|^{2}$ thus represents the probability that $|\psi\rangle$ is measured to be in the state $|\psi_{n}\rangle$. Given a hamiltonian $H$, or any other operator $A$ on ${\cal H}$, then $A^{\dagger}$ is defined in the standard way with respect to the above inner product.
(ii)   Modified metric for pseudo-hermitian Hamiltonians
Suppose the hamiltonian operator $H$ is pseudo-hermitian with the structure in (1), which we repeat here:
$H^{\dagger}={\cal K}\,H\,{\cal K}^{\dagger},~{}~{}~{}{\cal K}^{\dagger}{\cal K}=1,~{}~{}~{}~{}~{}{\cal K}^{\dagger}={\cal K},~{}~{}~{}~{}\Longrightarrow~{}~{}~{}{\cal K}^{2}=1.$ (26)
This leads us to define a new inner product which includes an insertion of the operator ${\cal K}$:
$\langle\!\langle\psi^{\prime}|\psi\rangle\!\rangle\equiv\langle\psi^{\prime}|{\cal K}|\psi\rangle.$ (27)
In simple words, kets have an accompanying ${\cal K}$ but bras do not, or visa-versa since ${\cal K}^{\dagger}={\cal K}$. In general, this new inner product has negative norm states. Below we will show that such states can be consistently projected out.
(iii)   Pseudo-hermitian operators correspond to observables with real eigenvalues
For any operator $A$ on ${\cal H}$, define its pseudo-hermitian conjugate as follows:
$A^{\dagger_{\scriptscriptstyle\kappa}}\equiv{\cal K}\,A^{\dagger}\,{\cal K}.$ (28)
Then the hamiltonian is pseudo-hermitian:
$H^{\dagger_{\scriptscriptstyle\kappa}}=H.$ (29)
$A^{\dagger_{\scriptscriptstyle\kappa}}$ is the proper conjugation based on the ${\cal K}$-metric in (27), namely
$\langle\!\langle\psi^{\prime}|A|\psi\rangle\!\rangle^{*}=\langle\!\langle\psi|A^{\dagger_{\scriptscriptstyle\kappa}}|\psi^{\prime}\rangle\!\rangle,$ (30)
where * denotes ordinary complex conjugation. We define a pseudo-hermitian operator $A$ as one that satisfies $A^{\dagger_{\scriptscriptstyle\kappa}}=A$. From (30), one concludes that pseudo-hermitian operators, in particular the hamiltonian H, has real eigenvalues. Any operator satisfying $A^{\dagger_{\scriptscriptstyle\kappa}}=A$ in principle can correspond to an observable. One can easily establish that this pseudo-hermitian adjoint satisfies the usual rules, e.g.
$(AB)^{\dagger_{\scriptscriptstyle\kappa}}$ $=$ $B^{\dagger_{\scriptscriptstyle\kappa}}A^{\dagger_{\scriptscriptstyle\kappa}}\,,$
$(aA+bB)^{\dagger_{\scriptscriptstyle\kappa}}$ $=$ $a^{*}A^{\dagger_{\scriptscriptstyle\kappa}}+b^{*}B^{\dagger_{\scriptscriptstyle\kappa}}\,,$ (31)
where $A,B$ are operators and $a,b$ are complex numbers.
(iv)   Conservation of probabilities
The hamiltonian determines the time evolution of a state $|\psi\rangle=|\psi(t=0)\rangle$, namely $|\psi(t)\rangle=U(t)\,|\psi\rangle$, where $U(t)=e^{-iHt}$. By conservation of probability we mean that the norm of states is preserved under time evolution. More generally:
$\langle\!\langle\psi^{\prime}(t)|\psi(t)\rangle\!\rangle=\langle\psi^{\prime}|e^{iH^{\dagger}t}\,{\cal K}\,e^{-iHt}\,|\psi\rangle=\langle\psi^{\prime}|{\cal K}e^{iH^{\dagger}t}{\cal K}^{2}\,e^{-iHt}\,|\psi\rangle=\langle\!\langle\psi^{\prime}(0)|\psi(0)\rangle\!\rangle$ (32)
In general, the modified ${\cal K}$-metric $\langle\!\langle\psi^{\prime}|\psi\rangle\!\rangle$ introduces negative norm states which imply negative probabilities. The above statement of unitarity just states that the sum of these probabilities is maintained under time evolution, but one must still deal with negative norm states which implies negative probabilities.
(v)   Projection onto positive norm states
For the main example of this paper, the ${\cal K}$-metric on the Hilbert space is indefinite, namely there are negative norm states which naively implies negative probabilities, which are clearly unphysical. However the properties of ${\cal K}$ naturally leads to projection operators which can fix this problem. Since ${\cal K}^{2}=1$, states can be classified according to ${\cal K}=\pm 1$. Namely,
${\cal K}|\psi_{n}\rangle=(-1)^{k_{n}}\,|\psi_{n}\rangle,~{}~{}~{}~{}k_{n}\in\{0,1\}.$ (33)
In fixing the norm of a state to be unity, the sum involves negative terms, which would be interpreted as negative probabilities:
$\langle\!\langle\psi|\psi\rangle\!\rangle=\sum_{n}(-1)^{k_{n}}\,|c_{n}|^{2}=1.$ (34)
Define the projectors
$P_{\pm}=\dfrac{1\pm{\cal K}}{2}.$ (35)
They satisfy the usual algebra of projectors:
$P_{+}+P_{-}=1,~{}~{}~{}~{}~{}P_{\pm}^{2}=P_{\pm},~{}~{}~{}~{}~{}P_{+}P_{-}=0.$ (36)
In addition, one has
${\cal K}P_{\pm}=P_{\pm}K=\pm P_{\pm}.$ (37)
Thus the original Hilbert space decomposes as
${\cal H}={\cal H}_{+}\oplus{\cal H}_{-},~{}~{}~{}~{}~{}|\psi_{\pm}\rangle\in{\cal H}_{\pm},~{}~{}~{}~{}~{}{\cal K}|\psi_{\pm}\rangle=\pm|\psi_{\pm}\rangle.$ (38)
One has
$\langle\!\langle\psi_{\pm}|\psi_{\pm}\rangle\!\rangle=\langle\psi|P_{\pm}{\cal K}P_{\pm}|\psi\rangle=\pm\langle\psi_{\pm}|\psi_{\pm}\rangle,~{}~{}~{}~{}~{}\langle\!\langle\psi_{-}|\psi_{+}\rangle\!\rangle=\langle\psi|P_{+}{\cal K}P_{-}|\psi\rangle=0$ (39)
If we project onto the sub Hilbert space ${\cal H}_{+}$, then all states $|\psi_{n}\rangle\in{\cal H}_{+}$ have positive norm, i.e. $k_{n}=0$, which implies positive probabilities $|c_{n}|^{2}\geq 0$. This positivity is preserved under time evolution:
$\langle\!\langle\psi_{\pm}(t)|\psi_{\pm}(t)\rangle\!\rangle=\langle\!\langle\psi_{\pm}(0)|\psi_{\pm}(0)\rangle\!\rangle,~{}~{}~{}~{}~{}\langle\!\langle\psi_{-}(t)|\psi_{+}(t)\rangle\!\rangle=\langle\!\langle\psi_{-}(0)|\psi_{+}(0)\rangle\!\rangle=0.$ (40)
In general there could be operators $A$ on ${\cal H}$ that mix ${\cal H}_{+}$ and ${\cal H}_{-}$. If one projects ${\cal H}$ onto ${\cal H}_{+}$, then one should restrict to operators which when acting on $|\psi_{+}\rangle$ yields a state that is also in ${\cal H}_{+}$. In other words
$\langle\!\langle\psi_{-}|A|\psi_{+}\rangle\!\rangle=0.$ (41)
It is easy to see that this requires $A$ to be ${\cal K}$-invariant:
$A={\cal K}A{\cal K}.$ (42)
The proof is simple:
$\langle\!\langle\psi_{-}|A|\psi_{+}\rangle\!\rangle=\langle\psi_{-}|{\cal K}A|\psi_{+}\rangle=-\langle\psi_{-}|A|\psi_{+}\rangle=\langle\psi_{-}|A{\cal K}|\psi_{+}\rangle=+\langle\psi_{-}|A|\psi_{+}\rangle,$ (43)
which implies (41). The condition (42) implies $A^{\dagger}={\cal K}A^{\dagger}{\cal K}$. Thus $A^{\dagger_{\scriptscriptstyle\kappa}}=A$ requires $A^{\dagger}=A$.
In summary, if we project onto the Hilbert space ${\cal H}_{+}$, then all probabilities are positive, and remain positive under time evolution. Let us now turn to the specifics of our model. The Hilbert space ${\cal H}$ has a basis consisting of n-particle states with specific momenta:
$|\psi\rangle=|({\bf{k}}_{1},s_{1}),({\bf{k}}_{2},s_{2}),\ldots({\bf{k}}_{n},s_{n})\rangle\equiv a^{\dagger}_{s_{1}}({\bf{k}}_{1})a^{\dagger}_{s_{2}}({\bf{k}}_{2})\cdots a^{\dagger}_{s_{n}}({\bf{k}}_{n})|0\rangle~{}~{}~{}~{}~{}s_{i}\in\{\pm\}.$ (44)
Mapping the labels $s_{i}=\pm$ to the integers $s_{i}=\pm 1$, define $\widehat{s}=(1-s)/2\in\{0,1\}$. Then
${\cal K}\,|\psi\rangle=(-1)^{\sum_{i}\widehat{s}_{i}}\,\,|\psi\rangle.$ (45)
This simple action on momentum space acts non-locally on the fields $\Phi(x)$ in Section II. In short, ${\cal K}=1$ for states with an even number of $a^{\dagger}_{-}({\bf{k}})$, whereas ${\cal K}=-1$ for an odd number. Projection onto states of positive norm with ${\cal K}=1$ leads to the Hilbert space ${\cal H}_{+}$ where states created by $a^{\dagger}_{-}$ must come in pairs, reminiscent of Cooper pairs.
Above we interpreted the operators $a_{\pm}^{\dagger}$ as creating ${\cal P}{\cal T}$ even/odd states. Thus projecting onto ${\cal H}_{+}$ breaks ${\cal P}{\cal T}$ symmetry.
IV Higher order RG beta-functions from the OPE
In this section we compute the beta functions to 3-loops for the fully anisotropic case $g_{1}\neq g_{2}\neq g_{3}$ based only on the OPE (18). We are more interested in the case where SU(2) is broken to U(1), where $g_{1}=g_{2}\neq g_{3}$ considered in the next section, however it’s easier to work out the beta functions for the fully anisotropic case first and then set $g_{1}=g_{2}$. Renormalization group beta-functions are generally prescription, or scheme, dependent. However for marginal perturbations, the one and two loop contributions are universal. [^5] To see this, consider a single coupling $g$ with $\beta_{g}=b_{2}g^{2}+b_{3}g^{3}+\ldots$ where $b_{2,3}$ are coefficients. The prescription dependence corresponds to a redefinition of the coupling $g^{\prime}=g^{\prime}(g)$. Let $g^{\prime}=g+cg^{2}+\ldots$. One easily sees that $\beta^{\prime}(g^{\prime})=b_{2}g^{\prime 2}+b_{3}g^{\prime 3}+\ldots.$ The implicit prescription of this article is based on the OPE (18), and doesn’t rely on any epsilon-expansion, nor Feynman diagrams for scalar fields. [^6] We do not address whether this OPE prescription can be shown to be equivalent to Feynman diagram perturbation theory and dimensional regularization. Referring to the title of Wilson-Fisher’s paper WilsonFisher , we work in precisely $4$ dimensions, not $3.99$!
We first review the 1-loop result in ALDolls , then describe how to extend this to higher orders. The starting point is the perturbative expansion about the conformal field theory for $\langle{\bf{X}}\rangle$ where ${\bf{X}}$ is an arbitrary operator (it can in fact to be taken to be the identity):
$\langle{\bf{X}}\rangle=\sum_{n=0}^{\infty}\dfrac{(-2\pi^{2})^{n}}{n!}\sum_{a_{i}}g_{a_{1}}g_{a_{2}}\cdots g_{a_{n}}\,\int d^{4}x_{1}d^{4}x_{2}\cdots d^{4}x_{n}\,\langle{\cal O}^{a_{1}}(x_{1}){\cal O}^{a_{2}}(x_{2})\cdots{\cal O}^{a_{n}}(x_{n})\,{\bf{X}}\rangle_{0},$ (46)
where $\langle X\rangle_{0}$ is the unperturbed CFT result. The $n$-th term gives a contribution to the $(n-1)$-loop $\beta$ function.
IV.1 1-loop
For the one-loop beta function one only needs the the OPE’s of the operators ${\cal O}^{a}$ themselves. Using (18) one finds
${\cal O}^{a}(x){\cal O}^{b}(y)=\dfrac{\delta^{ab}}{|x-y|^{8}}\,\,\dfrac{\kappa^{2}}{64\pi^{8}}-\dfrac{1}{16\pi^{4}|x-y|^{4}}\,\,C^{ab}_{c}\,\,{\cal O}^{c}(0)+\ldots$ (47)
where there is no sum over $c$, we used $\delta^{ab}f^{abc}=0$, and the only non-zero terms are
$C^{ab}_{c}=(f^{abc})^{2}=4~{}~{}~{}~{}~{}\forall~{}a\neq b\neq c.$ (48)
To order $g^{2}$:
$\langle{\bf{X}}\rangle=\langle{\bf{X}}\rangle_{0}-2\pi^{2}\sum_{c}g_{c}\int d^{4}x\,\langle{\cal O}^{c}(x)\,{\bf{X}}\rangle_{0}+2\pi^{4}\sum_{a,b}g_{a}g_{b}\int d^{4}x\int d^{4}y\,\langle{\cal O}^{a}(x)\,{\cal O}^{b}(y)\,{\bf{X}}\rangle_{0}+\dots\,,$ (49)
Using the OPE (47) along with
$\int_{a}\dfrac{d^{4}x}{x^{4}}=-2\pi^{2}\log a,$ (50)
where $a$ is an ultraviolet cut-off, one finds
$\langle{\bf{X}}\rangle=\langle{\bf{X}}\rangle_{0}-2\pi^{2}\sum_{a,b,c}\left(g_{c}-\dfrac{1}{8}C^{ab}_{c}g_{a}g_{b}\,\log a\right)\int d^{4}x\,\langle{\cal O}^{c}(x)\,{\bf{X}}\rangle_{0}+\dots\,.$ (51)
The ultraviolet divergence is removed by letting $g_{c}\to g_{c}+\tfrac{1}{8}C^{ab}_{c}g_{a}g_{b}\,\log a$. This leads to
$\beta_{g_{a}}\equiv\dfrac{dg_{a}}{d\ell}=\dfrac{1}{8}\sum_{b,c}\,C^{bc}_{a}\,g_{b}g_{c}$ (52)
where increasing $\ell=\log a$ is the flow to low energies. Thus to 1-loop we have
$\beta_{g_{1}}=g_{2}g_{3},~{}~{}~{}\beta_{g_{2}}=g_{1}g_{3},~{}~{}~{}\beta_{g_{3}}=g_{1}g_{2}.$ (53)
The above calculation indicates that our model is renormalizable to 1-loop, meaning that no additional operators, such as $J^{a}J^{a}$ or $\widetilde{J}^{a}\widetilde{J}^{a}$, are generated since the OPE (47) closes.
IV.2 2-loops
As we did above at 1-loop, at higher orders we use the OPE prescription to isolate $\log a$ divergences that contribute to the beta-functions. There are also $(\log a)^{n}$ divergences for $n>1$, which we interpret as arising from lower orders in perturbation theory which are already accounted for in the lower order beta-function. It turns out that the OPE (47) is insufficient to extract the appropriate $\log a$ contributions and one must split ${\cal O}^{a}$ into the product of $J^{a}$ and $\widetilde{J}^{a}$. It is useful to introduce diagrams that indicate the OPE structure that leads to a $\log a$ divergence of interest, with the building blocks shown in in Figure 1.777Unlike Feynman diagrams, these diagrams are not a prescription for calculating the full correlation function, but are useful for isolating $\log a$ divergences. The 1-loop contribution is shown in Figure 2.
Figure 1: OPE diagram building blocks. Each red dot signifies a spacetime point $x$. Solid lines refer to $J$ whereas $\widetilde{J}$ is represented by a dotted line.
Figure 2: 1-loop diagram
To order $g^{3}$ one has
$\langle{\bf{X}}\rangle_{{\rm 2-loop~{}term}}=-\dfrac{(2\pi^{2})^{3}}{3!}\sum_{a,b,c}g_{a}g_{b}g_{c}\,{\cal J}^{abc}$ (54)
where
${\cal J}^{abc}=\int d^{4}x_{1}\,d^{4}x_{2}\,d^{4}x_{3}\,\langle{\cal O}^{a}(x_{1}){\cal O}^{b}(x_{2}){\cal O}^{c}(x_{3}){\bf{X}}\rangle.$ (55)
Consider the contribution where ${\cal O}^{c}(x_{3})$ is left over after OPE’s:
${\cal J}^{abc}=-\dfrac{\kappa}{128\pi^{8}}\,\delta^{ab}\sum_{e}f^{ace}f^{bce}\int d^{4}x_{1}\,d^{4}x_{2}\,d^{4}x_{3}\,\dfrac{1}{|x_{1}-x_{2}|^{4}|x_{1}-x_{3}|^{2}|x_{2}-x_{3}|^{2}}\,\langle{\cal O}^{c}(x_{3}){\bf{X}}\rangle_{0}.$ (56)
This corresponds to the OPE diagram in Figure 3. One can perform the integral over $x_{1},x_{2}$ by shifting $x_{1}\to x_{1}+x_{3}$, $x_{2}\to x_{2}+x_{3}$, and using the integral:
$\int d^{4}x_{2}\,\dfrac{1}{|x_{1}-x_{2}|^{4}|x_{2}|^{2}}=\dfrac{\pi^{2}}{2}\dfrac{1}{|x_{1}|^{2}}.$ (57)
Then integrating over $x_{1}$ using (50) gives a $\log a$ divergence. This contribution comes with a combinatorial factor of $2\times 3!$, where the extra $2$ comes from the additional diagram that exchanges $J$ with $\widetilde{J}$. As for 1-loop we absorb this divergence in the couplings $g_{a}$. This leads to the beta functions
$\beta_{g_{1}}=g_{2}g_{3}-\dfrac{\kappa}{4}\,g_{1}(g_{2}^{2}+g_{3}^{2}),~{}~{}~{}~{}{\rm plus~{}cyclic~{}permutations~{}of~{}1,2,3}.$ (58)
Figure 3: The only 2-loop diagram that gives an additional contribution to the beta-functions.
The above, based on Figure 3, is the only contribution at order $g^{3}$. For instance the diagram in Figure 4 gives a $\log^{2}a$ divergence which is already incorporated in the 1-loop beta function. The diagram in Figure 5 gives zero since it is proportional to $\delta^{ab}f^{abc}$. Figure 6 diagram could potentially lead to new terms $J^{a}(x)J^{a}(x)$ in the action rendering our original model non-renormalizable, however as is evident from the diagram, the two $J$’s are at different spacetime points, and furthermore leads to $\log^{2}a$ divergences.
Figure 4: 2-loop diagram where the divergence is the product of two 1-loop divergences.
Figure 5: 2-loop diagram that gives zero contribution to the beta-function since it is proportional to $\delta^{ab}f^{abc}$.
Figure 6: 2-loop diagram that potentially gives rise to additional $JJ$ terms in the action.
IV.3 3 loops
There are 3 OPE diagrams that contribute to order $g^{4}$, shown in Figure 7. For each diagram the combinatorial factor is $2\cdot 4!$. The type A and C diagram contributions can be evaluated using the same integrals (50) and (57) as for 2-loops. This gives
$\beta_{g_{a}}^{\rm 3-loop\,A}=\beta_{g_{a}}^{\rm 3-loop\,C}=\dfrac{\kappa^{2}}{32}\sum_{b\neq c\neq a}g_{b}^{3}g_{c}~{}.$ (59)
The type B diagram is different since the OPE’s involving $f^{abc}$ are internal to the diagram. In this case one needs
$\int\dfrac{d^{4}y}{|x-y|^{4}|y|^{4}}=\dfrac{\pi^{2}}{|x|^{4}},~{}~{}~{}~{}~{}\int\dfrac{d^{4}y}{|x-y|^{2}|y|^{4}}=\dfrac{\pi^{2}}{4|x|^{2}}.$ (60)
This leads to
$\beta_{g_{a}}^{\rm 3-loop\,B}=\dfrac{\kappa^{2}}{32}\sum_{b\neq c\neq a}g_{a}^{2}g_{b}g_{c}~{}.$ (61)
Putting this all together, to 3-loops the beta functions are
$\beta_{g_{1}}=g_{2}g_{3}-\dfrac{\kappa}{4}g_{1}\left(g_{2}^{2}+g_{3}^{2}\right)+\dfrac{\kappa^{2}}{16}\left(g_{1}^{2}g_{2}g_{3}+g_{2}^{3}g_{3}+g_{3}^{3}g_{2}\right)+\ldots$ (62)
plus cyclic permutations of $1,2,3$.
Figure 7: 3-loop diagrams of type A, B and C.
V SU(2) broken to U(1) and higher orders
V.1 3 loops
The fully isotropic case $g=g_{1}=g_{2}=g_{3}$ has ${\rm SU(2)}$ symmetry since $\sum_{a}J^{a}\widetilde{J}^{a}$ is built upon the quadratic Casimir. Based on (62) the beta function is
$\beta_{g}=g^{2}-\dfrac{\kappa}{2}g^{3}+\dfrac{3\kappa^{2}}{16}g^{4}+\ldots$ (63)
The only fixed point is at $g=0$ which is marginally relevant or irrelevant depending on the sign of $g$. Recalling that our conventions for the RG flow are such that increasing $\ell$ corresponds to a flow to low energies, when $g>0$ one flows to $g=0$ in the UV where the theory is the asymptotically free theory of the $\Phi,\widetilde{\Phi}$ fields.
In this article we are more concerned with the above ${\rm SU(2)}$ symmetry broken to U(1). Setting $g_{1}=g_{2}\neq g_{3}$ in (62) one has
$\beta_{g_{1}}$ $=$ $g_{1}g_{3}-\dfrac{\kappa}{4}g_{1}(g_{1}^{2}+g_{3}^{2})+\dfrac{\kappa^{2}}{16}\,g_{1}^{2}(g_{3}^{3}+2g_{3}g_{1}^{2})+\ldots$ (64)
$\beta_{g_{3}}$ $=$ $g_{1}^{2}-\dfrac{\kappa}{2}\,\,g_{1}^{2}g_{3}+\dfrac{\kappa^{2}}{16}\,g_{1}^{2}(2g_{1}^{2}+g_{3}^{2})+\ldots$ (65)
For describing the RG trajectories in the next section, the analysis is greatly simplified by an RG invariant $Q(g_{1},g_{3})$, since this does not require integrating the beta functions as a function of scale $\ell$, since RG trajectories are constant $Q$ contours. Such an RG invariant $Q$ is defined as satisfying
$\sum_{g}\beta_{g}\partial_{g}Q=0.$ (66)
To 1-loop the RG invariant is $Q=g_{1}^{2}-g_{3}^{2}$. This RG invariant persists to 3-loops:
$Q=\left(g_{1}^{2}-g_{3}^{2}\right)\left(1+\dfrac{\kappa}{2}g_{3}+\dfrac{\kappa^{2}}{16}(3g_{3}^{2}+g_{1}^{2})\right).$ (67)
Here (66) is satisfied to order $g^{5}$, which implies it is valid to order $g^{4}$:
$\sum_{g=g_{1},g_{3}}\,\beta_{g}\partial_{g}Q={\cal O}(g^{5}).$ (68)
V.2 Conjecture for all orders
In going to higher orders, the OPE diagrams of the last section have nested geometric series which allows their resummation. To motivate this, note that for the fully isotropic case (63)
$\beta_{g}=g^{2}-\dfrac{\kappa}{2}g^{3}+\dfrac{3\kappa^{2}}{16}g^{4}+\ldots=\dfrac{g^{2}}{(1+g\kappa/4)^{2}}.$ (69)
In order to sum these higher order contributions, we compare with the current-current perturbations in 2D reviewed in the Appendix. Based on this we propose the following beta functions:
$\beta_{g_{1}}=\dfrac{g_{1}(g_{3}-g_{1}^{2}\kappa/4)}{(1-\kappa^{2}g_{1}^{2}/16)(1+\kappa g_{3}/4)},~{}~{}~{}~{}~{}\beta_{g_{3}}=\dfrac{g_{1}^{2}(1-\kappa g_{3}/4)^{2}}{(1-\kappa^{2}g_{1}^{2}/16)^{2}}.$ (70)
One can easily check that to order $g^{4}$, the above agree exactly with the 3-loop results (64). This is unexpected since the integrals involved are quite different. These beta functions preserve the RG invariant $Q$, which to all orders is now:
$Q=\dfrac{g_{1}^{2}-g_{3}^{2}}{(1-\kappa g_{3}/4)^{2}(1-g_{1}^{2}\kappa^{2}/16)}.$ (71)
The invariance condition (66) is exactly satisfied to all orders. The above beta functions do not constitute a gradient flow where $\beta_{g_{i}}=-\partial_{g_{i}}{\cal H}(g)$ for some height function ${\cal H}(g)$, since $\partial_{g_{1}}\beta_{g_{3}}\neq\partial_{g_{3}}\beta_{g_{1}}$. This is consistent with the existence of cyclic RG flows we find below since gradient flow is often associated with flows that are consistent with various forms of c-theorems.
The above beta functions and $Q$ have poles at the points $g_{1,3}=\pm 4$. These are self-dual points under the strong-weak coupling duality described below, equation (74), and the RG flows pass through them smoothly. The formulas (70) should be taken as a conjecture, since we did not analyze in full detail all the higher order OPE diagrams beyond 3-loops, however agreement with our 3-loop calculation is a non-trivial check. It should still be viewed as a conjecture since we did not prove that there are no additional contributions to the beta functions in a different prescription which are missed by our prescription based on the OPE. At worse, the above beta functions represent a re-summation of important contributions to the beta functions and can serve to provide some non-trivial checks of the RG flows we propose below.
VI Renormalization group trajectories
[…]
The beta functions (70) show that both $\beta_{g_{1}}$ and $\beta_{g_{3}}$ are zero when $g_{1}=0$, indicating a line of fixed points along the $g_{3}$ axis. The anomalous scaling dimension $\Gamma$ of the perturbation along this line is a function of $g_{3}$ and can be inferred from the slope of the beta function $\beta_{g_{1}}$. Generally speaking, in 4 spacetime dimensions, near a fixed point $g=g_{c}$
$\beta_{g}=(4-\Gamma)(g-g_{c}).$ (72)
From the leading term in the series for $\beta_{g_{1}}$ about $g_{1}=g_{c}=0$ one finds
$\Gamma(g_{3})=\dfrac{4}{1+g_{3}/4}.$ (73)
Based on this, there are 3 distinct regions along the $g_{3}$ line of fixed points:
(i) $g_{3}>0$. Here the perturbation is relevant with $40<\Gamma<$ 4.
(ii) $-4<g_{3}<0$. These are marginally irrelevant perturbations with $\Gamma>4$.
(iii) $g_{3}<-4$. These are still relevant perturbations but distinguished by $\Gamma<0$.
Figure 8: RG flows to one loop. Arrows indicate the flow to low energies.
It was shown in BLflow that RG flows based on the beta functions (70) can be extended to strong coupling $g_{1,3}\to\infty$ where the RG invariant $Q$ in (71) plays an essential role. This is due to a surprising strong-weak coupling duality
$g_{1,3}\to\widetilde{g}_{1,3}\equiv\dfrac{16}{g_{1,3}}.$ (74)
Namely, if $\widetilde{g}=16/g$, then
$\beta_{\widetilde{g}}(\widetilde{g})\equiv\dfrac{\partial\widetilde{g}}{dg}\,\beta_{g}=-\beta_{g}(g\to\widetilde{g}).$ (75)
The RG invariant $Q$ is also invariant under this strong-weak coupling duality:
$Q(\widetilde{g}_{1},\widetilde{g}_{3})=Q(g_{1},g_{3}).$ (76)
It was also shown in BLflow that it is consistent, and required, to endow the coupling constant space $(g_{1},g_{3})$ with the topology of a cylinder where one identifies $g_{3}=\pm\infty$, since $\beta(g_{1},g_{3})=\beta(g_{1},-g_{3})$ in the limit $|g_{3}|\to\infty$. This allows us to extend the flows to the entire $(g_{1},g_{3})$ plane, including through the poles at $g_{1}=\pm 4$, which are self-dual points under (74). We refer the reader to BLflow for more detailed explanations. This leads a variety of RG flows which we now itemize.
VI.1 Massive Flows with fixed points
For case (i) above, namely $g_{3}>0$, the perturbation is relevant $40<\Gamma<$ 4 and flows to strong coupling in the IR. We interpret this as a massive phase where the IR fixed point is empty. In Figure 9 we plot the constant $Q$ RG trajectories based only on the 3 loop result (67), and one sees that the 1-loop flowchart is not significantly modified. In comparison with the current-current perturbations reviewed in the Appendix, this is a sine-Gordon like phase.
Figure 9: Contour plot of the RG invariant $Q$ to 3-loops (67).
Next consider case (iii) above, namely $g_{3}<-4$. These are also relevant perturbations, with negative scaling dimension, which are also expected to be massive. $g_{3}=-4$ is a self-dual point, that is $\widetilde{g_{3}}=g_{3}$. One thus expects a different phase in this region. In comparison with 2D, this is analogous to a sinh-Gordon phase (see the Appendix).
A contour plot of the non-perturbative $Q$ in (71) indicates the RG flows shown in Figure 10. One sees that the set of all contours forms an interesting manifold. Flows in the cyclic regime are shown separately in Figure 12 since they are not clearly distinguishable in Figure 10. One see’s that they consistently cross the narrow bridges at the self-dual points.
Figure 10: Contour plot of the non-perturbative RG invariant $Q$ equation (71). The points $g_{3}=\pm\infty$ are identified.
VI.2 Massless Flows between two non-trivial fixed points for imaginary coupling
Massive theories typically terminate at an empty fixed point, since at low energies the massive particles decouple. On the other hand, flows that end in the IR at a non-trivial CFT are expected to be massless, since some massless degrees of freedom survive the flow. There are no such flows in our models if $g_{1}$ and $g_{3}$ are real. Well, since we are already considering non-unitary theories, consider $g_{1}$ to be an imaginary coupling:
$g_{1}\to ig_{1}.$ (77)
Then the resulting beta functions still have real coefficients and thus make sense. The RG invariant becomes
$Q=-\dfrac{g_{1}^{2}+g_{3}^{2}}{(1+g_{1}^{2}/16)(1-g_{3}/4)^{2}}.$ (78)
Thus at small coupling, the flows are no longer hyperbolas, but circles, and can thus can both originate and terminate along the line of fixed points. The trajectories based on the 3-loop $Q$ (67) are shown in Figure 11.
Figure 11: Contour Plot of the RG invariant $Q$ for $g_{1}$ imaginary. These flow contours begin and end along the critical $g_{3}$ line, and do not pass through it as the above contours do.
This implies non-trivial flows between two fixed points along the $g_{3}$ axis. These flows start from a relevant perturbation with dimension $\Gamma_{\scriptscriptstyle{\rm UV}}<4$ in the UV and arrive via an irrelevant perturbation in the IR with dimension $\Gamma_{\scriptscriptstyle{\rm IR}}>4$, as they must. An algebraic relation for these anomalous dimensions can be found using the RG invariant $Q$. When $g_{1}=0$, $Q=\left[g_{3}/(1-g_{3}/4)\right]^{2}$. Equating
$Q_{\scriptscriptstyle{\rm IR}}=Q_{\scriptscriptstyle{\rm UV}}~{}~{}~{}\Longrightarrow~{}~{}~{}g_{3}^{\scriptscriptstyle{\rm IR}}=\dfrac{2g_{3}^{\scriptscriptstyle{\rm UV}}}{g_{3}^{\scriptscriptstyle{\rm UV}}-2}.$ (79)
Using the relation (73), this can be expressed in terms of the $\Gamma$’s: [^8] At 1-loop, $g_{\scriptscriptstyle{\rm IR}}=-g_{\scriptscriptstyle{\rm UV}}$ and $\Gamma_{\scriptscriptstyle{\rm UV}}+\Gamma_{\scriptscriptstyle{\rm IR}}=2\cdot D=8$.
$\Gamma_{\scriptscriptstyle{\rm IR}}=\dfrac{3\Gamma_{\scriptscriptstyle{\rm UV}}-8}{\Gamma_{\scriptscriptstyle{\rm UV}}-3}.$ (80)
There is an interesting UV/IR duality, in that the above equation also implies
$\Gamma_{\scriptscriptstyle{\rm UV}}=\dfrac{3\Gamma_{\scriptscriptstyle{\rm IR}}-8}{\Gamma_{\scriptscriptstyle{\rm IR}}-3}.$ (81)
A consistency check of the above formula is that when $\Gamma_{\scriptscriptstyle{\rm UV}}=4$, then $\Gamma_{\scriptscriptstyle{\rm IR}}=4$, indicating no RG flow, which corresponds to the point $g_{1}=g_{3}=0$.
This regime of massless flows requires $\Gamma_{\scriptscriptstyle{\rm UV}}<4$ and $\Gamma_{\scriptscriptstyle{\rm IR}}>4$. Based on (80), this requires $\Gamma_{\scriptscriptstyle{\rm UV}}>3$. In terms of the coupling $g_{3}$ this corresponds to the regime
$33<\Gamma_{\scriptscriptstyle{\rm UV}}<4~{}~{}~{}~{}~{}\Longrightarrow~{}~{}~{}~{}0<g_{3}<4/$ 3 . (82)
One sees that $g_{3}=4/3$ is special point where $\Gamma_{\scriptscriptstyle{\rm IR}}=\infty$. We will say more about this in the next sub-section.
VI.3 Some special rational points in the 4D CFT’s
We have already seen that the point $g_{3}=4/3$ is special since it is at the boundary of the massless flows where $\Gamma_{\scriptscriptstyle{\rm IR}}=\infty$. One can surmise other potentially interesting special points as follows. When $\Gamma_{\scriptscriptstyle{\rm IR}}=\mathfrak{D}$ where $\mathfrak{D}$ is an integer $\mathfrak{D}\geq 4$, then the IR theory has some likely interpretations. For instance, if $\mathfrak{D}=8$, this could represent massless flows that arrive to the IR via the operator $T_{\mu\nu}T^{\mu\nu}$, or $(T_{\mu}^{\mu})^{2}$ where $T_{\mu\nu}$ is the dimension 4 energy-momentum tensor. If this were the case, then this is analogous to $T\overline{T}$ perturbations of 2D CFT SmirnovZamo . In this case $\Gamma_{\scriptscriptstyle{\rm UV}}=16/5$ which is rational and corresponds to $g_{3}=1$, which is in the range (82). If one repeats this argument for the same beta function in 2 spacetime dimensions based on (100), one finds $\Gamma_{\scriptscriptstyle{\rm UV}}=4/3$. This turns out to be the scaling dimension of the perturbation for the minimal ${\cal N}=2$ supersymmetric model at Virasoro central charge $c=1$ BLsusy , which is a special point of the sine-Gordon model at coupling $b^{2}=4/3$ in (96). Thus the 2D analog of this flow is such that it preserves the central charge $c$, but the ${\cal N}=2$ supersymmetry is broken in the flow and the remaining massless fields are goldstinos. [^9] To our knowledge this observation is absent from the literature. This flow is similar to massless flows that arrive in the IR to the CFT of a Majorana fermion via the dimension 4 operator $T\overline{T}$ Alyosha , where the UV central charge is $7/10$. It was shown in AhnLeClair that the only UV completions of flows that end at the Majorana description of the Ising model have UV central charge $c=7/10$ and $c=3/2$, both of which have ${\cal N}=1$ supersymmetry. [^10] There are other flows that arrive to the Ising model via $T\overline{T}$ based on a different spectrum related to the $E_{8}$ Lie algebra AhnLeClair , indicating multi-critical behavior. In the notation in Tanaka , this is a flow between $c<1$ minimal models $M(11,12)\to M(3,4)$, with central charge $c_{\scriptscriptstyle{\rm UV}}=21/22$, $c_{\scriptscriptstyle{\rm IR}}=1/2$. This flow is not included in Tanaka since there the focus on the generic SU(2) quantum group modular fusion category of Verlinde lines, whereas our $E_{8}$ based flow should be understood from the $E_{8}$ quantum group fusion category.
It is thus natural to consider massless flows that end with $\Gamma_{\scriptscriptstyle{\rm IR}}=\mathfrak{D}$ with $\mathfrak{D}$ an integer greater than $4$, since the IR fixed point involves irrelevant operators with integer dimensions and could correspond to Landau-Ginsburg theories where the scalar field $\phi$ has classical scaling dimension 1, or have other interpretations such as $T_{\mu\nu}T^{\mu\nu}$ in the IR. The result is quite simple:
$\Gamma_{\scriptscriptstyle{\rm IR}}=\mathfrak{D},~{}~{}~{}~{}\Longrightarrow~{}~{}~{}\Gamma_{\scriptscriptstyle{\rm UV}}=\dfrac{3\mathfrak{D}-8}{\mathfrak{D}-3},~{}~{}~{}~{}g_{3}^{\scriptscriptstyle{\rm UV}}=\dfrac{4(\mathfrak{D}-4)}{3\mathfrak{D}-8}.$ (83)
For $4\leq\mathfrak{D}\leq 8$:
$\{\Gamma_{\scriptscriptstyle{\rm UV}},~{}\mathfrak{D}=4,5,6,7,8\}$ $=$ $\{4,\tfrac{7}{2},\tfrac{10}{3},\tfrac{13}{4},\tfrac{16}{5}\}$ (84)
$\{g_{3}^{\scriptscriptstyle{\rm UV}},~{}\mathfrak{D}=4,5,6,7,8\}$ $=$ $\{0,\tfrac{4}{7},\tfrac{4}{5},\tfrac{12}{13},1\}.$ (85)
This points to the existence of rational CFT’s where important anomalous dimensions are rational numbers. The latter is of course contingent on the conjectured higher order beta functions (70).
The special point $g_{3}=4/3$ appears as the limit when $\mathfrak{D}=\infty$. From (83),
$\lim_{\mathfrak{D}\to\infty}g_{3}^{\scriptscriptstyle{\rm UV}}=\dfrac{4}{3}.$ (86)
In this limit $\Gamma_{\scriptscriptstyle{\rm UV}}=3$. This implies that the coupling $g_{1}$ has dimension $1$ in the UV, i.e. it is a like a mass coupling. One can argue that this special point is a free field theory, since there is nothing to flow to if $\Gamma_{\scriptscriptstyle{\rm IR}}>\infty$, thus there are effectively no flows. Much more work would be needed to establish this. However, in support of this idea, let us point out that for the models with the same beta function in 2D reviewed in the Appendix where (80) is replaced with (100), there the analog of this special point is $\Gamma_{\scriptscriptstyle{\rm UV}}=1$, which is known to be a mass term for a free Dirac fermion in 2D. [^11] Since relativistic fermion fields $\psi$ have dimension $3/2$ in 4D, this suggests that $g_{3}=4/3$ could also be a free fermion point, where the perturbation is $m\overline{\psi}\psi$ with $m$ identified by $g_{1}$. However lacking a formalism of bosonization in 4D, this would be difficult to establish at this stage.
VI.4 Cyclic RG flows
Let us return to our original model with $g_{1}$ real. When $g_{1}^{2}>g_{3}^{2}$, one is above the SU(2) symmetric separatrices and the flows never encounter the $g_{3}$ line of fixed points at $g_{1}=0$. These are the most exotic flows, since there are expectations that all QFT’s start and end at a fixed points, at least for unitary theories. We will return to addressing this below, but first we take the point of view that if our beta functions are correct, and if one is forced to understand all the flows in the $g_{1},g_{3}$ coupling constants, then cyclic flows in this regime are unavoidable.
In this regime $Q>0$. To 1-loop one can express the beta function $\beta_{g_{3}}$ in terms of $g_{3}$ and $Q$. Integrating this, the coupling constant as a function of the log of the length scale $\ell$ is
$g_{3}(\ell)=\sqrt{Q}\tan\left(\sqrt{Q}(\ell-\ell_{0})\right)$ (87)
where $\ell_{0}$ is an integration constant. The fundamental parameter of the theory is the period $\lambda$ of the RG, which is a simple function of $Q$ and thus an RG invariant:
$g_{3}(\ell+\lambda)=g_{3}(\ell),~{}~{}~{}~{}~{}\lambda=\dfrac{\pi}{\sqrt{Q}}.$ (88)
Thus one flows from $g_{3}=-\infty$ to $g_{3}=+\infty$ in a finite RG time $\lambda$. Note that along the SU(2) invariant separatrices, $Q=0$, such that the RG period goes to $\infty$.
This behavior could be spoiled at higher orders since the RG flows extend to $g_{3}=\pm\infty$ which is beyond the 1-loop weak coupling regime where (88) was derived. One can check that to 2-loops, the period of the cyclic RG remains as in (88). The all-orders beta functions (70) resolve this issue since these beta functions are well defined as $g\to\infty$. These non-perturbative expressions simply lead to a doubling of the RG period $\lambda$. To see this, we can eliminate $g_{1}$ from the beta function for $g_{3}$ using $Q$:
$\beta_{g_{3}}=\dfrac{dg_{3}}{d\ell}=16~{}\dfrac{\left(g_{3}^{2}-16Q(g_{3}-4)^{2}\right)\left(1-Q(g_{3}-4)^{2}\right)}{(g_{3}+4)^{2}}.$ (89)
The above can be integrated and one still finds a cyclic RG LeClairSierra , namely $g_{3}(\ell+\lambda)=g_{3}(\ell)$, where $\lambda$ is now twice the 1-loop result:
$\lambda=\int_{-\infty}^{\infty}\dfrac{dg_{3}}{\beta_{g_{3}}}=\dfrac{2\pi}{\sqrt{Q}}.$ (90)
In Figure 12 we plot the cyclic RG flows for a variety of positive $Q$ since they are difficult to isolate visually in Figure 10. One sees that the flows smoothly pass through the poles at the self-dual points $g_{1},g_{2}=\pm 4$. This can be attributed to the fact that the flows approach the self-dual points with the correct slope, namely along the SU(2) invariant flows along the diagonal. This can be seen in Figures 10, 12 where flows cross the poles through very narrow bridges.
Figure 12: Non-perturbative flows that are cyclic, based on the contour plot of the non-perturbative RG invariant $Q$ in equation (71) in the cyclic regime with $350.1<Q<$ 35. The points $g_{3}=\pm\infty$ are identified such that the topology of coupling constant space is that of a cylinder. Nowhere does the RG trajectory reach the fixed points along the $g_{3}$ axis.
Cyclic flows have previously been considered as rare, or just curiosities, because of various c-theorems and other considerationsctheorem ; CardyCtheorem ; Osborn ; JackOsborn ; Schwimmer ; Komargodski ; Polchinski . The latter studies often utilize quantum fields coupled to gravity since the focus is on the conformal, or trace anomaly of the stress-energy tensor. All of these works assume the theory is unitary. To address this further, first and foremost, our model certainly exists by construction, and as we showed, it has a regime with massless flows between conformal fixed points, which are more conventional flows. Although the CFT’s are non-unitary, these massless flows still fit in the paradigm that all flows begin and end at conformal fixed points. One should mention that in recent studies of massless flows between non-unitary minimal CFT’s in 2D, such flows are still consistent with a c-theorem Ravanini ; MussardoLG ; Tanaka ; KlebanovMM ; Negro , where here $c=c_{{\rm eff}}$, where $c_{{\rm eff}}$ is the effective central charge that appears in the thermodynamic Bethe ansatz. Returning to our model, if one is forced to understand the flows for all regimes of the couplings $g_{1},g_{3}$, then a cyclic regime where $Q$ is positive is an inescapable conclusion. At small coupling, this is just above the SU(2) symmetric diagonals. Secondly, for the 2D models with beta functions that are identical to those considered here and reviewed in the Appendix, an exact S-matrix was even proposed LeClairSierra , further justifying that these models are well-defined.
In K. Wilson’s pioneering work on the RG, he attempted to classify all possible RG flow behavior in a model independent way KWilson . Although he missed asymptotic freedom Wilczek ; Gross , he did point out the possibility of limit cycle behavior, and this was the subject of his last works with Glazek GW1 ; GW2 ; however the models in the latter work were not relativistic QFT’s. Returning to our problem, one needs to understand how our model circumvents certain, perhaps hidden assumptions behind studies that would seem to rule out cyclic RG flows. One obvious answer is that non-unitary theories, in particular those based on pseudo-hermitian hamiltonians, were not considered before. In Polchinski for instance, which relies heavily on the dilaton trick of Komargodski and Schwimmer Schwimmer ; Komargodski , the assumption of unitarity is explicitly stated. Besides this, it seems likely that in order to formulate something like a c-theorem, one needs a well-defined perturbation theory about both the UV and IR fixed points such that it applies to flows between fixed points, and this was also assumed in Polchinski . Clearly, the regime of our model with $Q>0$ has no such fixed points about which to perturb. For the 2D models reviewed in the Appendix, the “thermal central charge”, i.e. the one that is analyzed with the thermodynamic Bethe ansatz, indeed oscillates as a function of scale as one approaches the UV, and with a period that is correctly predicted by the beta functions Roman , namely $2\pi/\sqrt{Q}$. One should also mention that cyclic RG flows were found in a narrow, non-unitary region of couplings for $\phi^{6}$ theory in $D=3-\epsilon$ dimensions where $\phi$ is a matrix of scalar fields Klebanov . For a review of cyclic RG flows mainly aimed at applications to nuclear physics, see BraatenHammer .
VII Conclusions
We have computed the beta functions of the model in ALDolls to 3 loops, and showed that they do not spoil the 1-loop features found previously. For SU(2) broken to U(1) we re-summed an infinite number of contributions in order to obtain a non-perturbative beta function. This model has a line of fixed points, which are new non-unitary CFT’s in 4 spacetime dimensions. There exists RG flows between non-trivial fixed points and we computed the relation between the anomalous dimensions in the UV and IR, and furthermore identified some special points where these anomalous dimensions are rational. There also exists a regime where the flows do not begin nor end at a fixed point but in fact are cyclic, and we computed the period $\lambda$ of the flow in terms of the RG invariant $Q$. We argued that this circumvents the paradigm that all QFT’s begin or end at a fixed point as proposed in Schwimmer ; Polchinski , simply because the model is non-unitary.
The analysis of the higher order RG flows in this paper points to the existence of some new non-unitary CFT’s in 4 spacetime dimensions. We computed some anomalous dimensions of relevant perturbing operators in the UV and based on massless flows found some special rational exponents corresponding to flows that arrive in the IR via operators of integer dimension $\mathfrak{D}>4$. If these calculations are correct, they are still far from a complete understanding of these potential CFT’s: what are the primary fields and their correlation functions? It is interesting to observe that rational exponents can exist in both 2 and 4 spacetime dimensions, whereas in 3 dimensions, non-trivial exponents obtained by means of the epsilon expansion are highly non-perturbative and irrational WilsonFisher . This is perhaps due to there being no obvious analog to the $J^{a}\widetilde{J}^{a}$ factorization of marginal operators in $3D$, which would require $J^{a}$ to have dimension $3/2$ such as $\phi^{3}$, such that the marginal interaction goes as $\phi^{6}$. In 2D the obvious generalization of this factorization is as a product of left-right moving chiral conserved currents, which have dimension 1.
Our models may have applications to statistical mechanics in 4 spatial dimensions where there are less restrictive positivity constraints. However we argued that the pseudo-hermitian structure of the hamiltonian, namely $H^{\dagger}={\cal K}H{\cal K}$ with ${\cal K}^{2}=1$, allows a consistent projection onto positive norm states, and can perhaps still describe consistent unitary quantum mechanics for QFT in $3+1$ spacetime dimensions. The modified inner product based on the operator ${\cal K}$ is essential to this construction.
It would be clearly desirable to have a more fundamental understanding of the ${\cal K}$-insertions for our modified metric on the Hilbert space, since in this article it was just introduced in order to obtain the interesting OPE (18) since this structure leads to interesting beta-functions. Based on our results on massless flows above, by analogy with the classification of massless flows between non-unitary CFT;s in 2D Tanaka , we suggest that ${\cal K}$ can be interpreted as a topological defect, analogous to Verlinde lines in 2D CFT. In the 4D context, these should be codimension 2 defects, namely two dimensional surfaces. It is well accepted that such defects can modify the critical behavior in significant ways. The massless flows we found along the critical line $g_{1}=0$ should then be viewed as constrained by the U(1) fusion category. From this viewpoint, the line of fixed point we found are analogous to the radius of compactification of a free boson in 2D, which also has a duality. The flows in Tanaka are based on a study of the $SU(2)_{q}$ quantum group modular fusion category where $q$ is a root of unity, which is known to be a reduction of the U(1) category from various viewpoints. Very recent work of Witten is also suggestive of a such an understanding, where he argues that bras and kets can differ by an operator, which is analogous to our ${\cal K}$-inner product where kets have an operator ${\cal K}$ insertion, whereas bras do not.
There is some evidence for an underlying ${\rm SL}(2,\mathbb{Z})$ structure in our model. Above we showed that the higher order beta functions have a $g\to 1/g$ duality, like an S-duality $\tau\to-1/\tau$. T-duality, $\tau\to\tau+1$ typically corresponds to a periodicity due to a $\vartheta$-angle, as in ${\cal N}=4$ supersymmetric Yang-Mills theory. The cyclicity of the RG required us to identify $g_{3}=\pm\infty$, endowing the 2-parameter coupling constant space $(g_{1},g_{3})$ with the topology of a cylinder, which is a form of periodicity. This by itself does not establish an ${\rm SL}(2,\mathbb{Z})$ symmetry, since we have not precisely identified the elliptic modulus $\tau$ as a function of the 2 couplings. The role of ${\rm SL}(2,\mathbb{Z})$ is perhaps clearer for the fully anisotropic case with the beta functions (53) since integration of these 1-loop beta functions already leads to elliptic functions LeClairSierra . Furthermore the 2D models, with identical beta functions to our models, were proposed to have the S-matrix of Zamolodchikov Zelip built out of elliptic functions Elliptic . This S-matrix is characterized by a $\mathbb{Z}_{4}$ symmetry which is a subgroup of ${\rm SL}(2,\mathbb{Z})$, generated by the matrix $A=-i\sigma^{2}=\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}$. Furthermore, it is known that this S-matrix transforms covariantly under ${\rm SL}(2,\mathbb{Z})$, meaning the S-matrix transforms with phase factors or gauge transformations that preserve the structure of the Yang-Baxter equation Baxter ; Belavin . In Elliptic , for the SU(2) broken to U(1) model, which is the cyclic sine-Gordon model reviewed in the Appendix, the latter can be obtained from certain limits of Zamolodchikov’s elliptic S-matrix, and this is perhaps a way of understanding what is left of the ${\rm SL}(2,\mathbb{Z})$ structure for our two parameter model.
VIII Acknowledgements
We wish to thank Luis Alvarez-Gaumé, Denis Bernard, Igor Klebanov, Michael Peskin, Germán Sierra and Takahiro Tanaka for discussions.
IX Appendix: Comparison with 2D current-current perturbations
It is instructive to compare the above 3-loop results to exactly solvable models in 2 spacetime dimensions which have identical 1-loop beta functions, since this can provide insights into higher orders in perturbation theory, and more importantly, it provides a guide as to what the physical implications could be.
In Gerganov an all-orders beta function was proposed for the most general case of anisotropic current-current perturbations of Wess-Zumino-Witten models of 2D CFT for an arbitrary Lie group WittenWZW ; KZ . [^12] The paper Gerganov also included super Lie groups, the latter being motivated by applications to disordered systems and Anderson localization. The models are defined by the action
$S=S_{\rm WZW}+\sum_{A,a,b}g_{A}\,d^{A}_{ab}\,\int d^{2}x\,J^{a}(x)\overline{J}^{b}(x),$ (91)
where $J^{a}$, $\overline{J}^{a}$ are the left, right chiral conserved currents of the WZW model, $d^{A}_{ab}$ are fixed bi-linear coefficients, and $g_{A}$ are couplings. As in the present article, the prescription for the beta-functions was also based on the OPE of the currents $J^{a}$. In the conformal WZW model, the currents $J^{a},\overline{J}^{a}$ are functions of $z=x_{1}+ix_{2}$ and ${\overline{z}}=x_{1}-ix_{2}$ respectively, where $x_{1},x_{2}$ are the two euclidean spacetime coordinates. They satisfy the OPE
$J^{a}(z)J^{b}(0)=\dfrac{k\delta^{ab}}{2z^{2}}+\dfrac{1}{z}f^{abc}J^{c}(0)+\ldots$ (92)
and similarly for $\overline{J}({\overline{z}})$. This should be compared with the OPE (18). Above, $k$ is a fundamental parameter, the level of the affine Lie algebra, which is a positive integer for unitary theories. [^13] The conventions for the normalization of the 2D currents $J^{a}$ is different in minor ways, by factors of $\sqrt{2}$, from the definitions of the operators $J^{a}$ in the body of this paper. In this appendix, we do not pay close attention to overall factors, since for purposes of comparison with our 4D models, we normalize the couplings such that the 1-loop beta functions agree. For ${\rm SU(2)}$, the coefficients $d^{A}_{ab}$ can be chosen such that
$\sum_{A,a,b=1,2,3}g_{A}\,d^{A}_{ab}\,\,J^{a}\overline{J}^{a}=g_{1}\left(J^{+}{\overline{J}}^{-}+J^{-}{\overline{J}}^{+}\right)+g_{3}J^{3}\overline{J}^{3},$ (93)
where $J^{\pm}=J^{1}\pm iJ^{2}$. The beta-functions computed in Gerganov are precisely those in equation (70) with $\kappa=k$. In BLflow the RG invariant in (71) was found. For the SU(2) symmetric case $g_{1}=g_{3}$, the beta function agrees with the one proposed by Kutasov Kutasov .
When the level $k=1$, the currents can be bosonized in terms of the left/right components of a single scalar field $\phi=\varphi(z)+\overline{\varphi}({\overline{z}})$:
$J^{\pm}=\exp\left(\pm i\sqrt{2}\,\varphi\right),~{}~{}~{}~{}~{}J^{3}=i\partial_{z}\varphi.$ (94)
The action becomes
$S=\dfrac{1}{4\pi}\int d^{2}x\left(\tfrac{1}{2}(\partial\phi)^{2}+g_{1}\cos(\sqrt{2}\phi)+g_{3}(\partial\phi)^{2}\right).$ (95)
The $g_{3}$ term can be incorporated into the kinetic term, and by rescaling the scalar field, one obtains the sine-Gordon model:
$S=\dfrac{1}{4\pi}\int d^{2}x\left(\tfrac{1}{2}(\partial\phi)^{2}+g_{1}\cos(b\phi)\right),$ (96)
where $b$ is a function of the couplings $g_{1},g_{3}$ that can be found in BLflow . Henceforth we specialize to $k=1$.
Using the strong-weak coupling dualities (75) and (76) it was shown in BLflow that the flows can be completed to strong coupling. There are essentially 3 regions of couplings with different RG behavior depending on the value of $Q$. A sketch of these RG flows is shown in Figure 13; The actual trajectories can be determined from contour plots of $Q$ and are shown in Figures 10 and 12. The main features are the following:
$\bullet$ When $Q$ is negative with $|g_{3}|<4$, the coupling $b$ in (96) is real. This is a phase corresponding to marginally relevant or irrelevant perturbations of the sine-Gordon model.
$\bullet$ When $Q$ is negative with $|g_{3}|>4$, $b$ is imaginary. This phase corresponds to the sinh-Gordon model where $\cos(b\phi)$ becomes $\cosh(b\phi)$.
$\bullet$ When $Q$ is positive, the RG is cyclic, $g_{3}(\ell+\lambda)=g_{3}(\lambda)$, with period $\lambda=2\pi/\sqrt{Q}$. The S-matrix was proposed to be the analytic continuation of the usual sine-Gordon model ZamoZamo to the appropriate value of $b$ LeClairSierra . The resulting S-matrix has an infinite number of resonances with masses $m_{n}$ with Russian Doll scaling behavior
$m_{n}=2M_{s}\cosh(n\lambda/2)\approx M_{s}e^{n\lambda/2}~{}~{}~{}{\rm as~{}}n\to\infty$ (97)
where $n$ is an integer and $M_{s}$ the soliton mass LeClairSierra .
The anomalous dimensions $\Gamma(g_{3})$ along the critical line can be found from
$\beta_{g_{1}}=(2-\Gamma(g_{3}))g_{1}+\ldots.$ (98)
This gives
$\Gamma(g_{3})=\dfrac{2(4-g_{3})}{(4+g_{3})}.$ (99)
$\bullet$ When $g_{1}$ is imaginary, i.e. $g_{1}\to ig_{1}$, then this is the so-called imaginary sine-Gordon model Fendley1 ; Fendley2 , which has massless flows between two fixed points on the critical line $g_{1}=0$. These flows do not change the Virasoro central charge $c=1$, and the UV and IR fixed points are those of a free boson with different compactification radius. That the central charge does not change already violates the c-theorem ctheorem , presumably because of the non-unitarity. Quantum group reductions of these flows describe flows between unitary minimal models with $c<1$. Repeating the arguments of Section VI based on the beta functions, the relation between the anomalous dimensions in the UV and IR is
$\Gamma_{\scriptscriptstyle{\rm IR}}=\dfrac{\Gamma_{\scriptscriptstyle{\rm UV}}}{\Gamma_{\scriptscriptstyle{\rm UV}}-1}.$ (100)
These flows only exist for 1¡$\Gamma_{\scriptscriptstyle{\rm UV}}<2$, where $\Gamma_{\scriptscriptstyle{\rm UV}}=1$ corresponds to the free-fermion point of the sine-Gordon model. This model is integrable and the relation (100) agrees exactly with the thermodynamic Bethe-ansatz analysis in Fendley2 , and this provides a non-trivial check of the validity of our beta functions to all orders.
Figure 13: A rough sketch of non-perturbative flows for $g_{1}>0$. Flows for $g_{1}<0$ are a mirror reflection. A proper rendition of these flows based on the RG invariant $Q$ is shown in Figure 10. This figure borrowed from BLflow : sG, shG, and KT stand for sine-Gordon, sinh-Gordon, and Kosterlitz-Thouless.